If F(x) = x² + 2 and G(x) = 2/x, find (G ○ F)(2)
1
Expert's answer
2011-07-12T05:22:53-0400
If F(x) = x² + 2 and G(x) = 2/x, then

(G ○ F)(x) = 2/F(x) = 2/( x² + 2)
(G ○ F)(2) =& 2/( 2² + 2)& = 2/6 = 1/3